OpenStudy (anonymous):

Factor 45x2 - 36x A. 9(5x2 - 4x) B. 9x(5x -4) C. 36x(9 - x) D. None of the above

OpenStudy (primeralph):

B

OpenStudy (ajprincess):

45x^2=5*9*x*x -36x=-4*9*x what is common to both 45x^2 and -36x?

OpenStudy (ajprincess):

@nicc96

OpenStudy (anonymous):

@primeralph I request don't give answers directly.. For more details, check: http://openstudy.com/code-of-conduct

OpenStudy (skullpatrol):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.